October 8, 2014 Calibration
GlueX Calibration Meeting
Wednesday, October 8, 2014
11:00 am, EDT
JLab: CEBAF Center, F326
Connection Using Bluejeans
- To join via Polycom room system go to the IP Address: 184.108.40.206 (bjn.vc) and enter the meeting ID: 630804895.
- To join via a Web Browser, go to the page  https://bluejeans.com/630804895.
- To join via phone, use one of the following numbers and the Conference ID: 630804895
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
- More information on connecting to bluejeans is available.
- Status Update
- Commissioning Planning updates
- Fall 2014 Commissioning Simulations updates
- Data Monitoring
Attending: Sean, Simon, Mark I., Eugene, David. L, Mike S., Manuel, Eric, Kei, Will, Aristedes, other people in the counting room who I missed (JLab); Curtis, Paul, Naomi (CMU); Justin (MIT); Matt (IU)
We should have beam by next meeting!
- A few changes were made to CCDB. The start counter gain factor was changed to be the same as the BCAL to start out with, since they have similar readouts. New tables were added to store information on the target position and composition.
- New time-to-distance tables and B-field-dependent correction factors are being loaded as well. These are needed for the analysis of the fall commissioning data. However, they did cause degraded tracking efficiency in the simulation, since the simulation does not take its time-to-distance relation from the CCDB. Simon is working to update this.
- The new tables are loaded as different CCDB variations for purposes of testing. When data is taken, the correct tables will be associated with the correct runs.
- Naomi asked if the "no-field data" is going to be really with B=0. The current answer is yes. Probably a quench-safe field of 100A or 300A should make much difference at the initial precision we can reach.
Commissioning Planning Updates
- Beni has been running into successive problems taking data with various FDC crates. "It's like herding cats."
- For the FCAL, John. Z has successfully run the pi0 relative gain calibration code on simulated data generated using a pi0 particle gun. He applied a random gaussian variation to the cell energies and has been able to minimize the reconstructed pi0 width. The next step is to do the same with the commissioning simulation data, and using EVIO format data.
- Will McGinley has continued to look at the commissioning simulations for the BCAL. He looked at them with Sean and noticed a large number of hits on the upstream edge of the BCAL, presumably due to the many events generated upstream.
- The TOF group is planning to bootstrap their calibration using cosmic data that has events such as those with hits in two neighboring paddles.
- Sean has been helping Eric start up with Start Counter software in JANA. They found a "bug" in the legacy online monitoring plugin post-meeting.
- Plans for calibrating TDC output times without a reference to the accelerator RF clock are still developing.
- One easy way to bootstrap the TDC times would be to use the times from the fADCs on the same detector. This method requires the readout windows of the TDCs and fADCs to be aligned. Discussion in the online meeting that afternoon indicated that this is either current the case or is the plan to implement.
- The times can also be calibrated starting with the FCAL and working one's way upstream (e.g. FCAL -> TOF -> BCAL/SC -> PS -> Tagger).
- Sean asked about how easy it to get the results of the recent surveys. Mark pointed out that the data is pretty raw and needs a lot of analysis to get useful information out of. Eugene mentioned that currently there is a ~6mm shift between now and the last alignment, but it's not clear if this is due to systematic issues or what.
- Interalignment of detectors is soon going to be a serious issue. There are several standard software packages that could be used for global alignment which Sean has looked at. It would be useful to have someone look into this problem in greater detail.
- Simon fixed some start counter overlaps and added the new nose cone.
- Sean and David have been working on figuring out the cause of the job crashes due to bad BCAL hits. Not all of the data needed to diagnose the problem was being saved. Sean has provided some new files to David.
- After some discussion, it was agreed to evaluate producing more events on Friday.
- The current simulations are good enough for the detector groups to start analyzing and extracting rate information. The next steps agreed upon are:
- Detector groups should start looking at the simulations currently available to see if they make any sense and to extract information on expected rates.
- The output of the monitoring_hists plugin is a good place to start looking at. This output is available at /volatile/halld/detcom_01/hd_root
- Kei will try to run over the simulated EVIO files with the online plugins, which will give some more relevant information to look at.
- Mark and Sean will start running some EM background-only simulations.
- Kei reviewed the status of the online monitoring plugins that will be used by the online and offline monitoring efforts.
- Generally the CDC and BCAL are pretty well advanced, the other major detectors are fine, but the beamline subdetectors need more work. Notably, the pair spectrometer analysis code is missing from JANA.
- There are some weird things, like him not seeing DTAGHHits
- It would be useful to track the number of overflow and underflow values.
- Sean is in the process of updating the monitoring DB.
- Justin showed off some pretty sweet web pages, which allowed visualization of time series information, easy access to a bunch of different histograms, and detailed inspection of the contents of ROOT files through your web browser. You should check out the pages linked in the minutes.